Multiply:
(2x - 1)(2x2 - 3x + 2)

Answer by bandgeek1749(3)

(2x - 1)(2x2 - 3x + 2)
FOIL: first, outer, inner, last
first distrubute 2x:
2x X 2x2= 4x2
2x X -3x=-6x2
2x X 2= 4x
so this gives you equation 1: 4x2-6x2+4x
then you do the -1
-1 X 2x2= -2x2
-1 X -3x= 3x
-1 X 2= -2
so this gives you equation #2: -2x2 + 3x -2
you put the two equations together: 4x2-6x2+4x-2x2+3x-2
you add common terms so that gives you
-4x2-2x-2 as your answer :)